OpenStudy (anonymous):

Integrating Factors (equation soon to come)... Constants: alpha and k Initial Conditions: t=0, Cr=Cr*, Ca=Ca(0) 'naught' Final Conditions: t, Cr(t), Ca*

OpenStudy (anonymous):

\[\frac{ dC _{r} }{ dt }+\frac{ 1 }{ \alpha }C_r=0.5kC_a\]

OpenStudy (anonymous):

I think I am doing the integrating factor wrong

OpenStudy (anonymous):

\[e^{\int\limits_{?}^{?}\frac{ 1 }{ \alpha }dt}\]what are the limits? 0 to t ??

OpenStudy (anonymous):

$$\frac{dC_r}{dt}+\frac1\alpha C_r=\frac12kC_\alpha$$What exactly is \(C_a\)? What is \(r\)?

OpenStudy (anonymous):

two different unknowns. I need to solve for Cr(t) in my final answer. If you must know, Cr is the concentration of the reaction product and Ca is the concentration of the reactants

OpenStudy (anonymous):

So they are both constants here? Not functions of \(t\)?

OpenStudy (anonymous):

They are dependent on t

OpenStudy (anonymous):

if \[\frac{ dY(t) }{ dt }+P(t)Y=Q(t)\]Then\[Y(t)=Cr\]\[{1 \over \alpha}=P(t)\]and you get the pattern for Q(t)

OpenStudy (anonymous):

So is this a partial differentiation equation? If it's not,$$\frac{dC_r}{dt}+\frac1\alpha C_r=\frac12kC_a$$We want to multiply by some integration factor \(\mu\) so that our left-hand side is just the product rule, i.e.$$\mu\frac{dC_r}{dt}+\frac1\alpha\mu C_r=\frac{d}{dt}[\mu C_r]\\\mu\frac{dC_r}{dt}+\frac1\alpha\mu C_r=\mu\frac{dC_r}{dt}+\frac{d\mu}{dt}C_r$$Subtracting \(\mu\frac{dC_r}{dt}\) from both sides we find:$$\frac1\alpha\mu C_r=\frac{d\mu}{dt}C_r\\\frac1\alpha\mu=\frac{d\mu}{dt}$$... which is a separable differentiation equation in \(\mu\). Let's solve it by separating our variables:$$\frac1\alpha dt=\frac1\mu d\mu\\\int\frac1\alpha dt=\int\frac1\mu du\\\frac1\alpha t=\log\mu\\\mu=e^{\frac1\alpha t}$$Now we have our integration factor. Multiply throughout:$$e^{\frac1\alpha t}\frac{dC_r}{dt}+\frac1\alpha e^{\frac1\alpha t}C_r=\frac12k C_ae^{\frac1\alpha t}\\\frac{d}{dt}\left[e^{\frac1\alpha t}C_r\right]=\frac12kC_ae^{\frac1\alpha t}\\e^{\frac1\alpha t}C_r=\int\frac12kC_ae^{\frac1\alpha t} dt=\frac12kC_a\int e^{\frac1\alpha t} dt$$

OpenStudy (anonymous):

Evaluating that integral, we get $$e^{\frac1\alpha t}C_r=\frac12\alpha kC_ae^{\frac1\alpha t}+C\\C_r=\frac12\alpha kC_a+Ce^{-\frac1\alpha t}$$

OpenStudy (anonymous):

The solution is supposed to be.... which I am supposed to derive.

OpenStudy (anonymous):

The solution is what I found there. I just didn't use shortcut formulae for solving first-order linear equations like you wanted to. If that's what you want, consider that the solution for \(\dfrac{dy}{dt}+P(t)y=Q(t)\) can be written:$$y=\frac1\mu\int \mu Q(t)\,dt$$where$$\mu=e^{\int P(t)\,dt}$$ In our case, we have \(P(t)=\frac1\alpha\) and \(Q(t)=\frac12kC_a\), yielding $$\mu=e^{\int\frac1\alpha dt}=e^{\frac1\alpha t}$$and thus $$y=\frac1{e^{\frac1\alpha t}}\int\frac12kC_ae^{\frac1\alpha t}dt=e^{-\frac1\alpha t}\left(\frac12\alpha kC_ae^{\frac1\alpha t}+C\right)=\frac12\alpha kC_a+Ce^{-\frac1\alpha t}$$where \(C\) is our constant of integration.

OpenStudy (anonymous):

\[ C_r(t)=\frac{ 0.5k{\alpha} }{ 1+0.5k{\alpha} }C_{a_0}+(\frac{ C_{a_0} }{ 1+0.5k{\alpha} }-C_a^*)e^{-\frac{ (1+0.5k{\alpha})t }{ \alpha }}+(C_r^*+Ca^*-C_{a_0})e^{\frac{ -t }{ \alpha }}\]

OpenStudy (anonymous):

It's all there this time

OpenStudy (anonymous):

Thanks for helping, I have learned how to solve for the integrating factor. I'm going to use that and continue on. I think my chemical engineering teacher is breaking some rules but his notes are so hard to interpret. Thanks for your help

OpenStudy (anonymous):

No problem! I'm not sure where the other terms came from -- perhaps from solving for the initial conditions? Are there are other equations to solve?

OpenStudy (anonymous):

Ya he uses the initial and final conditions of Cr and Ca I posted in the header. These come into play when you integrate both sides, but I think this is assuming the integrating factor is treated as a constant.

OpenStudy (anonymous):

he shows this in his notes...

OpenStudy (anonymous):

for the left side

OpenStudy (anonymous):

\[\int\limits_{Y(t_0)}^{Y(t)} d[Y*\mu]=Y(t)\mu(t)-Y(t_0)\mu(t_0)\]